# Seepage analysis using a polygonal cell-based smoothed finite element method

### v1

@@ 4 Free-surface solutions
For steady seepage with a phreatic surface, the computational domain $\Omega$ is separated by an unknown free surface into a wetted region $\Omega_{w}$ and a dry region $\Omega_{d}$. The steady governing equation in the wetted region is written as
@@
Figure 1: Geometry and boundary conditions for the cubic seepage model.
Let the boundary of the seepage domain be decomposed as $\Gamma=\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}\cup\Gamma_{4}\cup\Gamma_{5}$, where $\Gamma_{1}$ is the upstream prescribed-head boundary, $\Gamma_{2}$ is the impervious boundary, $\Gamma_{3}$ is the unknown free surface, $\Gamma_{4}$ is the downstream seepage face, and $\Gamma_{5}$ is the downstream prescribed-head boundary corresponding to the tailwater level. The boundary conditions are given by
@@
where $\alpha$ is a small parameter and is taken as $\alpha=10^{-3}$ throughout this work. After solving for $h^{(t)}$, the pressure head is computed as $\psi^{(t)}(\mathbf{x})=h^{(t)}(\mathbf{x})-z$. The saturated region is then updated by
@@
The iteration is terminated when the change of the free-surface location becomes sufficiently small. In implementation, this can be monitored by the displacement of the downstream intersection (overflow) point $x_{o}$:
@@
Algorithm 1 Fixed-mesh free-surface iteration with an optional adaptive refinement loop
@@
3:  Build $k^{(0)}(\mathbf{x})$ from $\Omega_{w}^{(0)}$ using Eq.(33)
4:  for $t=0,1,\dots$ do $\triangleright$ inner fixed-mesh iteration
@@
6:   Solve Eq.(28) on the fixed mesh using the polygonal CSFEM to obtain $h^{(t)}$
7:   Compute $\psi^{(t)}=h^{(t)}-z$ and update $\Omega_{w}^{(t+1)}$ and the seepage face using Eq.(34)
8:   Update $k^{(t+1)}(\mathbf{x})$ using Eq.(33)
@@
14:  Mark elements in a narrow band around $\psi=0$ (or with large $\|\nabla h\|$) and locally refine the quadtree mesh with 2:1 balance
15:  Transfer the converged $h$ to the refined mesh by interpolation, set $\Omega_{w}^{(0)}$ from the transferred $\psi=h-z$, and continue
16:end for
The adaptive refinement step is optional and is particularly convenient in the present framework because hanging nodes created by local quadtree refinement are naturally handled by the hybrid quadtree–polygonal discretization without additional constraint equations or transition elements.
@@ 5 Numerical examples
$\mathbf{e}_{L_{2}}=\frac{\left|\mathbf{H}_{num}-\mathbf{H}_{ref}\right|_{L_{2}}}{\left|\mathbf{H}_{ref}\right|_{L_{2}}},$ (36)
@@ 5.1 Patch test
Table 1: Comparison of hydraulic head between the polygonal and quadtree meshes.
@@ 5.2.1 Foundation of a concrete dam
A convergence study is performed through mesh $h$-refinement with mesh sizes of $20~\mathrm{m}$, $10~\mathrm{m}$, $5~\mathrm{m}$, and $2.5~\mathrm{m}$. The mesh with an element size of $5~\mathrm{m}$ is shown in Fig. 5. Fig. 6 presents the relative errors of hydraulic head at the monitoring points, showing satisfactory convergence for both the polygonal CSFEM and the conventional FEM. Tab. 2 lists the hydraulic head values for the $5~\mathrm{m}$ mesh. The relative errors for the CSFEM and FEM are $7.1\times 10^{-3}$ and $1.1\times 10^{-2}$, respectively, indicating that the CSFEM achieves higher computational accuracy. Furthermore, Fig. 7 shows that the hydraulic head contours obtained by the CSFEM are in excellent agreement with those of the FEM.
@@ 5.3.1 Dam foundation with irregular geometry
In this example, a concrete dam with an irregular foundation is analyzed, as illustrated in Fig. 10. To investigate the transient seepage behavior, two monitoring locations, C and D, are chosen within the foundation domain. The storage coefficient is taken as $S_{s}=1.0\times 10^{-3}~\mathrm{m^{-1}}$, and the hydraulic conductivity is $k=1.0\times 10^{-3}~\mathrm{m/min}$. The initial upstream and downstream water levels are 0.0 m and 0.0 m, respectively. Over time, the upstream hydraulic head is progressively increased from 0.0 m to 2.0 m, and eventually reaches 4.0 m, as illustrated in Fig. 11.
@@
Figure 13: Hydraulic head boundary condition in the upstream.
@@ 5.3.2 Transient seepage analysis for complex geometry
Fig. 15 presents the temporal evolution of the hydraulic head at the monitoring points. The results obtained using the CSFEM with polygonal elements and hybrid quadtree elements exhibit excellent agreement with the reference solution. Moreover, Fig.16 shows the hydraulic head distribution at different time. Similarly, the hydraulic head distributions obtained by the CSFEM with polygonal elements and hybrid quadtree elements match the reference solution very well.
@@ 5.4.1 Free-surface seepage in a homogeneous dam
The benchmark considers a homogeneous rectangular dam, whose geometry is shown in Fig. 17. The dam has a height of 1.0 m and a base width of 0.5 m. A constant hydraulic head of 1.0 m is prescribed along the upstream boundary, whereas the downstream boundary is maintained at a head of 0.5 m. The foundation boundary is assumed to be impermeable. To adequately capture the free-surface zone forming above the downstream water level, two a priori locally refined meshes with the same refinement pattern are employed, as shown in Fig. 18a and b. Specifically, the potential phreatic-zone region is pre-refined to a cell size of 0.0125 m, whereas the remainder of the domain adopts a coarser size of 0.05 m. The conforming polygonal mesh is shown in Fig. 18a, while the corresponding quadtree mesh is shown in Fig. 18b. In addition, an adaptive quadtree mesh is constructed via solution-driven local refinement in the vicinity of the predicted free-surface region, as illustrated in Fig. 18c, so that resolution is concentrated where steep hydraulic gradients and active-set switching occur while keeping the total degrees of freedom moderate.
@@
Figure 18: Mesh model of a homogeneous dam; (a) the polygonal mesh; (b) the quadtree mesh; (c) the adaptive mesh.
@@ 5.4.2 Homogeneous trapezoidal dam
Figure 22: Mesh model of a trapezoidal dam; (a) the polygonal mesh; (b) the quadtree mesh; (c) the adaptive mesh.
@@ 6 Conclusions
This paper developed a polygonal CSFEM framework for two-dimensional seepage analyses covering steady-state, transient, and free-surface problems. Wachspress interpolation was employed on polygonal elements, and a cell-based gradient smoothing strategy was adopted so that the element matrices can be assembled using boundary integrals only. Based on the numerical investigations presented in this work, the following conclusions can be drawn.
(1) The proposed formulation successfully passes the patch test on both polygonal and quadtree discretizations, reproducing the linear hydraulic-head field with relative errors on the order of $10^{-8}$, which confirms the linear completeness and numerical stability of the smoothed formulation on different mesh topologies.
(2) For steady and transient seepage benchmarks involving regular and irregular geometries, including domains with curved boundaries and complex cavities, the polygonal CSFEM solutions agree well with FEM reference results and available analytical or literature solutions, demonstrating good accuracy and robustness for practical seepage simulations.
(3) For free-surface seepage in homogeneous dams, the fixed-mesh iterative strategy combined with the polygonal CSFEM yields free-surface profiles and pressure-head contours that are smooth and physically consistent, and the computed overflow-point locations show relative errors on the order of $10^{-2}$ across the investigated mesh discretizations. The quadtree and adaptive meshes tend to produce slightly smoother free-surface curves than the polygonal mesh, benefiting from a more regular element layout and locally enhanced resolution.
(4) The computational studies confirm the efficiency advantages of quadtree-based refinement and adaptivity within the present framework. Compared with the polygonal discretization, the quadtree mesh substantially reduces degrees of freedom and runtime while preserving accuracy, and the adaptive mesh further achieves comparable solution quality with markedly reduced mesh size and CPU time by concentrating refinement near the phreatic surface and other high-gradient regions.
Future work will focus on extending the present framework to heterogeneous and anisotropic conductivity fields, more rigorous a posteriori error indicators for adaptive refinement, and coupled hydro-mechanical applications where evolving seepage fields interact with deformation and stability.

### v2

@@ 4 Free-surface solutions
For steady seepage with a phreatic surface, the computational domain $\Omega$ is separated by an unknown free surface into a wetted region $\Omega_{w}$ and a dry region $\Omega_{d}$. For simplicity, an isotropic conductivity $k$ is assumed in the free-surface examples below. The steady governing equation in the wetted region is written as
@@
Figure 2: Geometry and boundary conditions for the free-surface seepage model.
Let the boundary of the seepage domain be decomposed as $\Gamma=\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}\cup\Gamma_{4}\cup\Gamma_{5}$, where $\Gamma_{1}$ is the upstream prescribed-head boundary, $\Gamma_{2}$ is the impervious boundary, $\Gamma_{3}$ is the unknown free surface, $\Gamma_{4}$ is the downstream seepage face, and $\Gamma_{5}$ is the downstream prescribed-head boundary representing the tailwater level. The boundary conditions are given by
@@
where $\alpha$ is a small parameter and is taken as $\alpha=10^{-3}$ throughout this work. After solving for $h^{(t)}$, the pressure head is computed as $\psi^{(t)}(\mathbf{x})=h^{(t)}(\mathbf{x})-z$.
The saturated region is then updated by
@@
The iteration is terminated when the change of the free-surface location becomes sufficiently small. In implementation, this can be monitored by the displacement of the downstream intersection point $x_{o}$:
@@
4.1 Adaptive refinement indicator
The adaptive refinement step is optional and aims to improve resolution around the active-set switching region and steep hydraulic gradients. Let $K$ denote an element in the current mesh $\mathcal{T}_{h}$. We first define a free-surface band indicator
$I_{K}=\begin{cases}1,&\min\limits_{\mathbf{x}\in\mathcal{S}(K)}\psi(\mathbf{x})\leq 0\leq\max\limits_{\mathbf{x}\in\mathcal{S}(K)}\psi(\mathbf{x}),\\ 0,&\text{otherwise},\end{cases}$ (41)
where $\mathcal{S}(K)$ denotes a set of sampling points on $K$. This indicator detects elements intersected by the zero level set $\psi=0$.
To prioritize refinement within this band, a gradient indicator is further defined as
$G_{K}=\left\|\nabla h\right\|_{0,K}=\left(\int_{K}\left|\nabla h\right|^{2}\,\mathrm{d}\Omega\right)^{1/2},$ (42)
where $\nabla h$ is evaluated from the smoothed hydraulic-gradient operator on $K$. The refinement indicator is then taken as
$\eta_{K}=I_{K}\,G_{K},$ (43)
so that refinement is restricted to a narrow band around the free surface while emphasizing elements with large hydraulic gradients.
4.2 Marking strategy and stopping criterion
Given $\eta_{K}$ on all $K\in\mathcal{T}_{h}$, a bulk-chasing (Dörfler) marking strategy is adopted [10, 33]: we select a minimal set of elements $\mathcal{M}\subset\mathcal{T}_{h}$ such that
$\sum_{K\in\mathcal{M}}\eta_{K}^{2}\geq\theta\sum_{K\in\mathcal{T}_{h}}\eta_{K}^{2},$ (44)
with $\theta\in(0,1)$, and refine all elements in $\mathcal{M}$ while enforcing a 2:1 balance in the quadtree mesh. The optional adaptivity loop is terminated when the fixed-mesh iteration satisfies Eq. (40) and the global indicator norm
$\|\eta\|_{\ell^{2}(\mathcal{T}_{h})}=\left(\sum_{K\in\mathcal{T}_{h}}\eta_{K}^{2}\right)^{1/2}$ (45)
falls below a prescribed tolerance. In practice, the adaptivity loop may also be stopped when the marked set $\mathcal{M}$ becomes empty.
Algorithm 1 Fixed-mesh free-surface iteration with optional adaptivity
@@ 4.2 Marking strategy and stopping criterion
3:  Build $\mathbf{k}^{(0)}(\mathbf{x})$ from $\Omega_{w}^{(0)}$ using Eq. (38)
4:  for $m=0,1,\dots$ do $\triangleright$ inner fixed-mesh iteration
@@
6:   Solve Eq. (33) on the fixed mesh using the polygonal CSFEM to obtain $h^{(m)}$
7:   Compute $\psi^{(m)}=h^{(m)}-z$ and update $\Omega_{w}^{(m+1)}$ and the seepage face using Eq. (39)
8:   Update $\mathbf{k}^{(m+1)}(\mathbf{x})$ using Eq. (38)
@@
14:  Compute $\eta_{K}$ using Eqs. (41)–(43) and mark elements by Eq. (44)
15:  if $\|\eta\|_{\ell^{2}(\mathcal{T}_{h})}$ in Eq. (45) is below the prescribed tolerance or $\mathcal{M}=\emptyset$ then
16:   stop and return the converged solution
17:  end if
18:  Locally refine the quadtree mesh with 2:1 balance and rebuild hybrid polygonal cells around hanging nodes
19:  Transfer the converged $h$ to the refined mesh by interpolation, set $\Omega_{w}^{(0)}$ from the transferred $\psi=h-z$, and continue
20:end for
The adaptive refinement step is particularly convenient in the present framework because hanging nodes created by local quadtree refinement are naturally handled by the hybrid quadtree discretization without additional constraint equations or transition elements.
@@ 5 Numerical examples
$e_{L_{2}}=\frac{\left\|\mathbf{H}_{num}-\mathbf{H}_{ref}\right\|_{L_{2}}}{\left\|\mathbf{H}_{ref}\right\|_{L_{2}}},$ (46)
@@ 5.1 Patch test
Table 1: Comparison of relative error between the polygonal and quadtree meshes.
@@ 5.2.1 Foundation of a concrete dam
A convergence study is performed through mesh $h$-refinement with mesh sizes of $20~\mathrm{m}$, $10~\mathrm{m}$, $5~\mathrm{m}$, and $2.5~\mathrm{m}$. The mesh with an element size of $5~\mathrm{m}$ is shown in Fig. 6. Fig. 7 presents the relative errors of hydraulic head at the monitoring points, indicating that the polygonal CSFEM and the conventional FEM exhibit comparable convergence rates under mesh refinement. Tab. 2 lists the hydraulic head values for the $5~\mathrm{m}$ mesh. The relative errors for the CSFEM and FEM are $7.1\times 10^{-3}$ and $1.1\times 10^{-2}$, respectively, indicating that the CSFEM achieves higher computational accuracy. Furthermore, Fig. 8 shows that the hydraulic head contours obtained by the CSFEM are in excellent agreement with those of the FEM.
@@ 5.3.1 Dam foundation with irregular geometry
In this example, a concrete dam with an irregular foundation is analyzed, as illustrated in Fig. 11. To investigate the transient seepage behavior, two monitoring locations, C and D, are chosen within the foundation domain. The storage coefficient is taken as $S_{s}=1.0\times 10^{-3}~\mathrm{m^{-1}}$, and the hydraulic conductivity is $k=1.0\times 10^{-3}~\mathrm{m/min}$. The initial upstream and downstream water levels are 0.0 m and 0.0 m, respectively. Over time, the upstream hydraulic head is progressively increased from 0.0 m to 2.0 m, and eventually reaches 4.0 m, as illustrated in Fig. 12. A reference solution is obtained using a sufficiently refined FEM mesh.
Fig. 13 shows the transient hydraulic head responses at monitoring points C and D, where the CSFEM predictions are in close agreement with the reference solution and the conventional FEM, capturing the gradual head rise induced by the staged upstream boundary condition. Fig. 14 presents the hydraulic head contours at the final time, demonstrating a smooth distribution consistent with the FEM reference and confirming the robustness of the proposed CSFEM on irregular dam-foundation geometry.
@@
Figure 14: Hydraulic head distribution in the dam foundation.
@@ 5.3.2 Transient seepage analysis for complex geometry
Fig. 16 presents the temporal evolution of the hydraulic head at the monitoring points. The results obtained using the CSFEM with polygonal elements and hybrid quadtree elements exhibit excellent agreement with the reference solution. Moreover, Fig. 17 shows the hydraulic head distribution at different times. Similarly, the hydraulic head distributions obtained by the CSFEM with polygonal elements and hybrid quadtree elements match the reference solution very well.
@@ 5.4.1 Free-surface seepage in a homogeneous dam
The benchmark considers a homogeneous rectangular dam, whose geometry is shown in Fig. 18. The dam has a height of 1.0 m and a base width of 0.5 m. A constant hydraulic head of 1.0 m is prescribed along the upstream boundary, whereas the downstream boundary is maintained at a head of 0.5 m. The foundation boundary is assumed to be impermeable. To adequately capture the free-surface zone forming above the downstream water level, two a priori locally refined meshes with the same refinement pattern are employed, as shown in Fig. 19(a) and (b). Specifically, the potential phreatic-zone region is pre-refined to a cell size of 0.0125 m, whereas the remainder of the domain adopts a coarser size of 0.05 m. The conforming polygonal mesh is shown in Fig. 19(a), while the corresponding quadtree mesh is shown in Fig. 19(b). Moreover, an adaptive quadtree mesh is constructed via solution-driven local refinement in the vicinity of the predicted free-surface region, as illustrated in Fig. 19(c), so that resolution is concentrated where steep hydraulic gradients and active-set switching occur while keeping the total degrees of freedom moderate.
@@
Figure 19: Mesh model of a homogeneous dam; (a) the polygonal mesh; (b) the quadtree mesh; (c) the adaptive mesh. The adaptive mesh in (c) is the final mesh obtained by the indicator in Eqs. (41)–(42) with bulk-chasing marking (Eq. (44), $\theta=0.5$) and 2:1 balanced quadtree refinement.
@@ 5.4.2 Homogeneous trapezoidal dam
Figure 23: Mesh model of a trapezoidal dam; (a) the polygonal mesh; (b) the quadtree mesh; (c) the adaptive mesh. The adaptive mesh in (c) is the final mesh obtained by the indicator in Eqs. (41)–(42) with bulk-chasing marking (Eq. (44), $\theta=0.5$) and 2:1 balanced quadtree refinement.
@@ 6 Conclusions
A polygonal CSFEM framework has been developed for two-dimensional seepage analyses covering steady-state, transient, and free-surface problems. By combining Wachspress interpolation with cell-based gradient smoothing, element matrices are assembled via boundary integrals only, avoiding in-element derivatives and improving robustness on distorted and locally refined meshes. Based on the numerical results presented in this work, the following conclusions are drawn.
(1) The proposed formulation passes the patch test on both polygonal and quadtree meshes, reproducing linear hydraulic-head fields with relative errors on the order of $10^{-8}$, which verifies linear completeness and numerical stability on different mesh topologies.
(2) For steady-state and transient seepage benchmarks on regular and irregular domains, including curved boundaries and complex cavities, the polygonal CSFEM solutions agree closely with FEM reference results and available analytical or literature solutions, demonstrating accuracy and robustness for practical seepage simulations. In a representative steady seepage example, the CSFEM yields smaller hydraulic-head errors than a conventional FEM at the same characteristic mesh size.
(3) For free-surface seepage in homogeneous dams, the fixed-mesh iteration coupled with the polygonal CSFEM produces smooth free-surface profiles and physically consistent pressure-head contours. Across the investigated discretizations, the computed overflow-point locations exhibit relative errors on the order of $10^{-2}$, and quadtree-based meshes generally yield slightly smoother free-surface curves due to their more regular layout and locally enhanced resolution.
(4) Quadtree refinement and solution-driven adaptivity significantly improve computational efficiency within the proposed framework. Compared with polygonal discretization, quadtree meshes reduce the degrees of freedom and runtime while preserving accuracy, and adaptive meshes further achieve comparable solution quality with markedly reduced mesh size and CPU time by concentrating refinement near the phreatic surface and other high-gradient regions.
Future work will extend the present framework to heterogeneous and anisotropic conductivity fields, develop more rigorous a posteriori error indicators for adaptive refinement, and investigate coupled hydro-mechanical applications in which evolving seepage fields interact with deformation and stability.